Coefficient of friction for a rough incline is 0.50 . find the acceleration of a block down the incline of angle 2A where A is the angle of repose

Grade:11

1 Answers

vikas askiitian expert
13 years ago

u=tanA =0.5

a=gsin2A-ugcos2A

  =g(2tanA-u(1-tan^2A))/1+tan^2A                          sin2A=2tanA/1+tan^2A

   a=g/2                                                                          cos2A=1-tan^2A/1+tan^2A